OpenStudy (anonymous):

can someone show me the steps to solve (6x^-2)^2(0.5x)^4 I know the answer is 2.25

OpenStudy (anonymous):

If you don't know the steps to this equation, how did you solve it?

OpenStudy (phi):

to "solve" you need an equation (an = sign) or to get a number, you need a value for x

OpenStudy (anonymous):

1 (((6x - 2) • 2) • (— • x)) • 4 2

OpenStudy (anonymous):

3.1 Pull out like factors : 6x - 2 = 2 • (3x - 1)

OpenStudy (anonymous):

x (4 • (3x - 1) • —) • 4 2

OpenStudy (anonymous):

2x • (3x - 1) • 4

OpenStudy (jdoe0001):

as phi said, there's no " = " sign, thus, there's nothing to solve

OpenStudy (phi):

is the question to simplify \[ (6x^{-2})^2(0.5x)^4 \]?

OpenStudy (phi):

you should learn a few "rules" first x^2 (for example) is short-hand for x*x if we did \( (x^3)^2 \) that means \(x^3 \cdot x^3 \) also \( x^3 = x\cdot x \cdot x\) so \[ (x^3)^2 = x^3 \cdot x^3 \\= x\cdot x \cdot x \cdot x\cdot x \cdot x \] and that is hard to write, so we use the short hand, and write \[ x^6\] now notice this "quick way": in \( (x^3)^2 \) 2*3 is 6 it turns out \( (x^3)^2 = x^{2\cdot 3}= x^6 \)

OpenStudy (phi):

also if you see (a*b)^2 that means (a*b)*(a*b) which can also be written as a*a*b*b (we can switch the order when multiplying) a*a*b*b in short hand is \(a^2 b^2\) the idea is \[ (a\cdot b)^2 = a^2 b^2 \] we can use these ideas to simplify the first part of your problem \[ (6x^{-2})^2 = 6^2 (x^{-2})^2 \] 6*6 is 36 what is \( (x^{-2})^2\) ? use the rule up above: it is \[x^{-2\cdot 2}\] which is \(x^{-4} \) so the first part simplifies to \[ (6x^{-2})^2 =36 x^{-4}\]

OpenStudy (phi):

can you write out (0.5x)^4 ?